OpenStudy (anonymous):

Is the equation -2 +- sqrt20 in its simplest form

OpenStudy (raden):

u just need simplify of radical form only

OpenStudy (raden):

sqrt(20) = sqrt(4*5) = sqrt(4) * sqrt(5) = 2 sqrt(5)

OpenStudy (anonymous):

thanks

OpenStudy (raden):

welcome :)

OpenStudy (anonymous):

anybody need an answer